Ashley is 10 years older than Brandon. In seven years, Ashley will be twice as old as Brandon. How old are they?
Found 2 solutions by stanbon, addingup:
Answer by stanbon(75887)

Ashley is 10 years older than Brandon. In seven years, Ashley will be twice as old as Brandon. How old are they?
------
A = B + 10
A+7 = 2(B+7)
------
Substitute for "A" and solve for "B"::
B+10 + 7 = 2B + 14
-----
B = 17-14 = 3 (Brandon's age now)
----
A = B + 10 = 13 (Ashley's age now)
--------------
Cheers,
Stan H.
------------

Answer by addingup(3677)

A = B+10
A+7 = 2(B+7)
B+10+7 = 2(
B+17 = 2B+14
B+3 = 2B
3 = B
Brandon is 3 and Ashley is 3+10 = 13.
In 7 years Ashley will be 13+7=20
and Brandon: 3+7 = 10. So Ashley is 2x10 = 20 just like the problem says.